OpenStudy (anonymous):

Divide. Write each quotient in simplest form. 3/5 ÷ 1/5 = _____ the answer is 3/25 is that right?

OpenStudy (vishweshshrimali5):

no it is 3

OpenStudy (jamesj):

No. Think intuitively for a moment. Suppose you slice a pizza up into 5 equal slices. So a slice is 1/5. If you have 3/5; i.e. 3 slices, how many 1/5 do you have? That's what 3/5 ÷ 1/5 is asking.

OpenStudy (anonymous):

3

OpenStudy (turingtest):

remember dividing by a fraction is the same as multiplying by it's reciprocal: (3/5)/(1/5)=(3/5)*(5/1)=3 as usual JamesJ's explanation is interesting though

OpenStudy (anonymous):

thank you i understand how to do it now:)